A food packet is dropped from the plane going at an altitude of 100m. What is the path of the packet as seen from the plane? What is the path of the packet as seen from the ground? If someone asks “what is the actual path”, what will you answer?

Grade:12

3 Answers

bibhash jha
13 years ago

The path of the packet seen from the plane will be a free fall in straight verticle line .

The path seen from the ground wil be a parabola . The  actual path can only be defined from a particular frame of reference .

Tarun
5 years ago
Let velocity if plane=v=initial velocity if of packet .so wrt ground it wiĺl be parabola if v is not zero and it will be free fall wrt plane it is moving with constant velocity v
Akshit
4 years ago
U=2m/s
T=2s
a=-9.8m/square because it is released downwards 
So using v=u+at 
V=(2)+(2)(-9.8)
V=2-19.6 that is v=17.6 m/s